Find the area of the region in the XY-plane enclosed by y = 3−x and x = 3y−y . In doing so, sketch the region (hint: remember that the graph of a quadratic is a parabola), and be sure to show all your work.
Answer #1

\\\mathrm{solution}\\ \\ \mathrm{y=3-x,\:x=3y-y}\\ \\ \mathrm{A=\int _a^b|f\left(x\right)-g\left(x\right)|dx}\\ \\ \mathrm{x=3y-y}\\ \\ \mathrm{3y-y=x}\\ \\ \mathrm{add \;similar\;elements;3y-y=2y}\\ \\ \mathrm{2y=x}\\ \\ \mathrm{\mathrm{Divide\:both\:sides\:by\:}2}\\ \\ \mathrm{\frac{2y}{2}=\frac{x}{2}}\\ \\ \mathrm{\frac{2y}{2}=\frac{x}{2}}\\ \\ \mathrm{y=\frac{x}{2}}\\ \\ \mathrm{f\left(x\right)=3-x}\\ \\ \mathrm{g\left(x\right)=\frac{x}{2}}\\ \\

\\\mathrm{a=2,\:b=3}\\ \\ \mathrm{=\int _2^3\left|3-x-\frac{x}{2}\right|dx}\\ \\ \mathrm{\mathrm{Find\:the\:equivalent\:expressions\:to\:}\left|3-x-\frac{x}{2}\right|\mathrm{\:at\:}2\le \:x\le \:3\mathrm{\:without\:the\:absolutes}}\\ \\ \mathrm{=\int _2^3\left(\frac{3x}{2}-3\right)dx}\\ \\ \mathrm{=\int _2^3\frac{3x}{2}dx-\int _2^33dx}\\ \\ \mathrm{\Rightarrow \int _2^3\frac{3x}{2}dx=\frac{3}{2}\cdot \int _2^3xdx=\frac{3}{2}\left[\frac{x^{1+1}}{1+1}\right]_2^3=\frac{3}{2}\left[\frac{x^2}{2}\right]_2^3=\frac{5}{2}=\frac{3}{2}\cdot \frac{5}{2}=\frac{15}{4}}\\ \\ \mathrm{\Rightarrow \int _2^33dx=\left[3x\right]_2^3=3}\\ \\ \mathrm{=\frac{15}{4}-3}\\ \\ \mathrm{=\frac{3}{4}}\\ \\ \mathrm{\mathrm{The\:area\:is:}}\\ \\ \mathrm{=\frac{3}{4}}\\ \\ \mathrm{graph \;is \;below}